oration has submitted the following sales forecast for the upcoming fiscal year (all sales are d account): 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter 13,580 Budgeted unit sales 11,580 12,588 14,5ee The selling price of the companys product is $14 per unit. Managerment expects to collect 75% of sales in the quarter in which the sales are made, 20% in the following quarter, and 5% of sales are expected to be uncollectible. The beginning balance of accounts receivable, all of which is expected to be collected in the first quarter, is $71,200 The company expects to start the first quarter with 1,725 units in finished goods inventory. Management desires an ending finished goods inventory in each quarter equal to 15% of the next quarters budgeted sales. The desired ending finished goods inventory for the fourth quarter is 1,925 units. Required: 1. Calculate the estimated sales for each quarter of the fiscal year and for the year as a whole 2. Calculate the expected cash collections for each quarter of the fiscal year and for the year as a whole 3. Calculate the required production in units of finished goods for each quarter of the fiscal year and for the year as a whole Complete this question by entering your answers in the tabs below
Required: 1. Calculate the estimated sales for each quarter of the fiscal year and for the year as a whole. 2. Calculate the expected cash collections for each quarter of the fiscal year and for the year as a whole. 3. Calculate the required production in units of finished goods for each quarter of the fiscal year and for the year as a whole Complete this question by entering your answers in the tabs below 39 Required 1 Required 2 Required 3 Calculate the estimated sales for each quarter of the fiscal year and for the year as a whole. 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Year Total sales Required 2>
Required: 1. Calculate the estimated sales for each quarter of the fiscal year and for the year as a whole 2. Calculate the expected cash collections for each quarter of the fiscal year and for the year as a whole 3. Calculate the required production in units of finished goods for each quarter of the fiscal year and for the year as a whole Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Calculate the expected cash collections for each quarter of the fiscal year and for the year as a whole. 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Year Total cash collections KRequired 1 Required 3>
Required: 1. Calculate the estimated sales for each quarter of the fiscal year and for the year as a whole 2. Calculate the expected cash collections for each quarter of the fiscal year and for the year as a whole 3. Calculate the required production in units of finished goods for each quarter of the fiscal year and for the year as a whole Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required3 Calculate the required production in units of finished goods for each quarter of the fiscal year and for the year as a whole. 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Year Required production in units Required 2 Reqaired 3
Answer #1

Answer:-1)-

Jessi Corporation
Estimated Sales
For the Year ended
Particulars 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Total
Budgeted sales units (a) 11500 12500 14500 13500 52000
Selling price per unit (b) $14 $14 $14 $14 $14
Estimated sales $ (c=a*b) 161000 175000 203000 189000 728000

2)-

Jessi Corporation
Schedule of Collections from sales
For the Year ended
Particulars 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Total
$ $ $ $ $
Opening receivable (a) 71200 0 0 0 71200
Cash Collection (b)
75% in quarter of sale 161000*75%=120750 175000*75%=131250 203000*75%=152250 189000*75% =141750 546000
20% in following quarter of sale 161000*20%=32200 175000*20%=35000 203000*20%=40600 107800
Total cash collections (c=a+b) 191950 163450 187250 182350 725000

3)-

Jessi Corporation
Calculation of production units
For the Year ended
Particulars 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Total
Budgeted units sales 11500 12500 14500 13500 52000
Add:- Closing invetory of finished goods
15% of the next quarter's budgeted sales 12500*15%=1875 14500*15%=2175 13500*15%=2025 1925 8000
Less:-Opening invetory of finished goods 1725 1875 2175 2025 7800
No. of units to be produced 11650 12800 14350 13400 52200